Gallium has 2 natural occuring isotopes with mass number of 69 and 71 repectively. What is the percentage if both their relatively atomic mass 70.59amu?
1
Expert's answer
2021-07-30T05:37:02-0400

let the percentage of gallium 69 =x

percentage of gallium 71=100-x

mass of gallium 71= 70.59

let the mass of gallium 69=68.9257

atomic mass of gallium from the periodic table = 69.723

"69.723=\frac{68.9257x +70.59(100-x)}{100}"

"6972.3=68.9257x-70.59x+7059"

"1.6643x=86.7"

"x=52.09\%"

percentage of gallium 61=52.09"\%"

percentage of gallium 71 = 100-52.09=47.91"\%"